# Statistical Test on Diffusion Model-based Generated Images by Selective Inference

### v2

@@ 4.1 Conditional Distribution of Test Statistics
Due to the complexity described in the previous section, it is difficult to directly obtain the sampling distribution of $T(\bm{X})$. Then, we consider the sampling distribution of $T(\bm{X})$ conditional on the event that the anomaly region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomaly region $\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$, i.e.,
$T(\bm{X})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (11)
In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameters that is independent of the test statistic. As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space. The nuisance parameter $\mathcal{Q}_{\bm{X}}$ is written as
$\mathcal{Q}_{\bm{X}}=\left(I_{n}-\frac{\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}}\bm{\nu}_{\mathcal{M}_{\bm{X}}}^{\top}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}}}\right)\bm{X}.$ (12)
@@
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X})|>|T(\bm{X}^{\mathrm{obs}})|\mid\bm{X}\in\mathcal{X}\right),$ (13)
@@
$\mathcal{X}=\left\{\bm{X}\in\mathbb{R}^{n}\mid\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}},\mathcal{Q}_{\bm{X}}=\mathcal{Q}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (14)
@@
$\mathcal{X}=\left\{\bm{X}(z)\in\mathbb{R}^{n}\mid\bm{X}(z)=\bm{a}+\bm{b}z,z\in\mathcal{Z}\right\},$ (15)
where vectors $\bm{a},\bm{b}\in\mathbb{R}^{n}$ are defined as
$\bm{a}=\mathcal{Q}_{\bm{X}^{\mathrm{obs}}},\;\bm{b}=\frac{\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}},$ (16)
@@
$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}\right\}.$ (17)
Let us consider a random variable $Z\in\mathbb{R}$ and its observation $z^{\mathrm{obs}}\in\mathbb{R}$ so that they satisfy $\bm{X}=\bm{a}+\bm{b}Z$ and $\bm{X}^{\mathrm{obs}}=\bm{a}+\bm{b}z^{\mathrm{obs}}$. Then, the selective $p$-value in (13) is re-written as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right).$ (18)
Under the null hypothesis $\mathrm{H}_{0}$, the distribution of the unconditional variable $Z$ is $\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\Sigma\bm{\nu}_{\mathcal{M}_{\bm{X}}})$. Consequently, given $Z\in\mathcal{Z}$, the conditional random variable $Z$ adheres to a truncated Gaussian distribution. Once the truncated region $\mathcal{Z}$ is identified, computing the selective $p$-value in (18) becomes straightforward. Therefore, the remaining task is the identification of $\mathcal{Z}$.
@@ 4.2 Over-Conditioning
To compute the truncated region $\mathcal{Z}$, we employ a divide and conquer approach. It is difficult to directly identify the truncated region $\mathcal{Z}$ due to the complexity of the computational algorithm of the diffusion model. The basic idea of this approach is to decompose the data space $\mathcal{X}$ into a set of polyhedra by considering additional conditioning, which we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022). It is easy to understood that a polyhedron in the $n$-dimensional data space ${\mathcal{X}}$ corresponds to an interval in the one-dimensional space ${\mathcal{Z}}$. Therefore, we can sequentially examine intervals in the one-dimensional space and check whether the same hypothesis (anomaly region) as the observed one is selected. In this study, we show that the filtered reconstruction error $E(\bm{X})$ can be expressed as a piecewise-linear function of $\bm{X}$. By exploiting this, we identify a over-conditioned interval $\mathcal{Z}^{\mathrm{oc}}\subset\mathcal{Z}$.
@@ 4.3 Identification of $\mathcal{Z}$ by Parametric Programming
$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}}\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z).$ (24)
The number of $\mathcal{Z}^{\mathrm{oc}}$ is obviously finite due to the finiteness of the number of polyhedra, but for practical purposes it grows exponentially, making it difficult to identify all of them. In many other SI studies, it is known that a search from $z_{\mathrm{min}}=(-10\sigma-|z^{\mathrm{obs}}|)$ to $z_{\mathrm{max}}(=10\sigma+|z^{\mathrm{obs}}|)$ is sufficient for practical use, where $\sigma$ is the standard deviation of the test statistic $T(\bm{X})$. An algorithm for calculating the selective $p$-value via Parametric Programming is summarized in Algorithm 2.
@@
0:  $\mathbf{X^{\mathrm{obs}}},z_{\mathrm{min}},z_{\mathrm{max}}$ and $z^{\mathrm{obs}}:=T(\bm{X}^{\mathrm{obs}})$
@@
2:  Obtain $\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$ by (7)
@@
6:     Compute $\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})$ and $\mathcal{M}_{\bm{a}+\bm{b}z}$ by (23) for $z$
7:     if $\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{X}^{\mathrm{obs}}}$ then
8:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})$
@@
10:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b})}+\varepsilon$, where $\varepsilon$ is small positive number.
@@
12:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z^{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right)$
@@ 5.1 Numerical Experiments
Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$. The details of the structure of the diffusion model we used are shown in Appendix C. The diffusion models were trained on data tailored to each experimental settings. The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=200$, and the reconstruction was conducted $5$ step samplings. The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear. In this experiment, we aim to generate new images through probabilistic sampling, $\eta$ was set to $1$. We used only normal images for the type I error rate experiments. The synthetic dataset for normal images is generated to follow $\bm{X}=(X_{1},X_{2},\dots,X_{n})\sim\mathcal{N}(\bm{0},\Sigma)$. we made 1000 normal images for $n\in\left\{64,256,1024,4096\right\}$. In the power experiments, we used only anomaly images. We made 1000 anomaly images $\bm{X}=(X_{1},X_{2},\dots,X_{n})\sim\mathcal{N}(\bm{\mu},\Sigma)$, where $\mu_{i}=\Delta,\forall i\in\mathcal{S}$, and $\mu_{i}=0,\forall i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomaly region with its position randomly chosen. The image size of the anomaly images was set to 256, with signals $\Delta\in\left\{1,2,3,4\right\}$. The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $2$. All experiments were conducted under the significance level $\alpha=0.05$. In addition, we conducted robustness experiments against non-Gaussian noise, i.e., $\bm{X}\not\sim\mathcal{N}$. The details are described in Appendix D.
@@ Results.
Figure 2 shows type I error rate. The proposed methods DMAD-test and DMAD-test-oc can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$. In contrast, naive and permutation cannot control the type I error rate. Figure 3 shows powers. Since naive and permutation cannot control the type I error rate, their powers are not considered. Among the methods that can control the type I error rate, the proposed method has the highest power. DMAD-test-oc is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power.
@@ 5.2 Real Data Experiments
We used the dataset in (Buda et al., 2019). This dataset includes brain images with tumors as normal images (939 images) and brain images without tumors as anomaly images (941 images). A random selection of 100 images from the normal images was made to estimate the mean and variance. The remaining images were standardized using the estimated mean and variance, and 750 normal images were extracted for training the diffusion model. The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=330$, with reconstruction performed through 5 step samplings. We set the threshold $\lambda=0.8$ and the kernel size of the averaging filter to $2$. The results of the DMAD-test and naive are shown in the Figure 4. The left column represents the results for normal brain images without tumors, while the right column represents the results for abnormal brain images with tumors. The $p_{\rm selective}$ calculated by the proposed DMAD-test is high for normal images (True Negative) and low for anomaly images (True Positive), indicating that the results are desirable. On the other hand, the $p_{\rm naive}$ obtained by the naive test is low not only for anomaly images but also for normal images (False Positive), indicating the invalidness of the naive test. This result indicates that the DMAD-test detected anomaly region as statistically significant while avoiding misidentification of normal image as anomaly.
$p_{\rm naive}=0.027$, $p_{\rm selective}=0.866$ Refer to caption $p_{\rm naive}=0.028$, $p_{\rm selective}=0.611$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.530$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.784$ Refer to caption $p_{\rm naive}=0.001$, $p_{\rm selective}=0.937$ Refer to caption $p_{\rm naive}=0.009$, $p_{\rm selective}=0.145$
(a) Brain images without tumor
$p_{\rm naive}=0.000$, $p_{\rm selective}=0.000$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.032$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.001$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.006$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.000$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.036$
(b) Brain images with tumor
@@
Figure 4: Example results of the proposed DMAD-test and the naive method applied to brain images.
6 Conclusions, Limitations and Future Works
In this study, we proposed a statistical test for anomaly regions in medical images detected by a diffusion model. With the proposed DMAD-test, the false detection probability can be controlled with the desired significance level because statistical inference is properly conducted conditional on the fact that the anomaly regions are identified by a diffusion model. We demonstrated the proposed method on a simple anomaly detection task, but its applicability to other tasks remains unclear. Moreover, as the complexity of the diffusion model architecture increases, so does the computational cost. Therefore, to apply the proposed method to practical problems, it is necessary to improve computational efficiency.

### v3

@@ 4.1 Conditional Distribution of Test Statistics
Due to the complexity described in the previous section, it is difficult to directly obtain the sampling distribution of $T(\bm{X},\bm{X}^{\mathrm{ref}})$. Then, we consider the sampling distribution of $T(\bm{X},\bm{X}^{\mathrm{ref}})$ conditional on the event that the anomalous region $\mathcal{M}_{\bm{X}}$ is the same as the observed anomalous region $\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$, i.e.,
$T(\bm{X},\bm{X}^{\mathrm{ref}})\mid\left\{\mathcal{M}_{\bm{X}}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}\right\}.$ (14)
In the context of SI, to make the characterization of the conditional sampling distribution manageable, we also incorporate conditioning on the nuisance parameter that is independent of the test statistic. As a result, the calculation of the conditional sampling distribution in SI can be reduced to a one-dimensional search problem in an $n$-dimensional data space. The nuisance parameter $\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}$ is written as
$\mathcal{Q}_{\bm{X},\bm{X}^{\mathrm{ref}}}=\left(I_{2n}-\frac{\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}\bm{\nu}_{\mathcal{M}_{\bm{X}}}^{\top}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}}}\right)\begin{pmatrix}\bm{X}\\ \bm{X}^{\mathrm{ref}}\end{pmatrix}.$ (15)
@@
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{X},\bm{X}^{\mathrm{ref}})|>|T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})|\mid\bm{X}\in\mathcal{X}\right),$ (16)
@@
[…]
@@
$\mathcal{X}=\left\{\begin{pmatrix}\bm{X}(z)\\ \bm{X}^{\mathrm{ref}}(z)\end{pmatrix}\in\mathbb{R}^{2n}\middle|\begin{pmatrix}\bm{X}(z)\\ \bm{X}^{\mathrm{ref}}(z)\end{pmatrix}=\bm{a}+\bm{b}z,z\in\mathcal{Z}\right\},$ (18)
where $\bm{X}(z)=\bm{a}_{1:n}+\bm{b}_{1:n}z$, and $\bm{c}_{1:n}$ represents a vector composed of the first $n$ elements of the vector $\bm{c}$. The vectors $\bm{a},\bm{b}\in\mathbb{R}^{2n}$ are defined as
$\bm{a}=\mathcal{Q}_{\bm{X}_{\mathrm{obs}}},\;\bm{b}=\frac{\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}}{\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}},$ (19)
@@
$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}\right\}.$ (20)
Let us consider a random variable $Z\in\mathbb{R}$ and its observation $z_{\mathrm{obs}}\in\mathbb{R}$ so that they satisfy $\bm{X}=\bm{a}_{1:n}+\bm{b}_{1:n}Z$ and $\bm{X}_{\mathrm{obs}}=\bm{a}_{1:n}+\bm{b}_{1:n}z_{\mathrm{obs}}$. Then, the selective $p$-value in (16) is re-written as
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right).$ (21)
Under the null hypothesis $\mathrm{H}_{0}$, the distribution of the unconditional variable $Z$ is $\mathcal{N}(0,\bm{\nu}^{\top}_{\mathcal{M}_{\bm{X}}}\tilde{\Sigma}\bm{\nu}_{\mathcal{M}_{\bm{X}}})$. Consequently, given $Z\in\mathcal{Z}$, the conditional random variable $Z$ adheres to a truncated Gaussian distribution. Once the truncated region $\mathcal{Z}$ is identified, computing the selective $p$-value in (21) becomes straightforward. Therefore, the remaining task is the identification of $\mathcal{Z}$.
@@ 4.2 Over-Conditioning
To compute the truncated region $\mathcal{Z}$, we employ a divide and conquer approach. It is difficult to directly identify the truncated region $\mathcal{Z}$ due to the complexity of the computational algorithm of the diffusion model. The basic idea of this approach is to decompose the data space $\mathcal{X}$ into a set of polyhedra by considering additional conditioning, which we refer to as over-conditioning (OC) (Duy and Takeuchi, 2022). It is easy to understand that a polyhedron in the $n$-dimensional data space ${\mathcal{X}}$ corresponds to an interval in the one-dimensional space ${\mathcal{Z}}$. Therefore, we can sequentially examine intervals in the one-dimensional space and check whether the same hypothesis (anomalous region) as the observed one is selected. In this study, we show that the filtered reconstruction error $E(\bm{X})$ can be expressed as a piecewise-linear function of $\bm{X}$. By exploiting this, we identify a over-conditioned interval $\mathcal{Z}^{\mathrm{oc}}\subset\mathcal{Z}$.
@@ 4.3 Identification of $\mathcal{Z}$ by Parametric Programming
$\mathcal{Z}=\bigcup_{z\in\mathbb{R}\mid\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}}\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z).$ (27)
The number of $\mathcal{Z}^{\mathrm{oc}}$ is obviously finite due to the finiteness of the number of polyhedra, but for practical purposes it grows exponentially, making it difficult to identify all of them. In many other SI studies, it is known that a search from $z_{\mathrm{min}}=(-10\sigma-|z_{\mathrm{obs}}|)$ to $z_{\mathrm{max}}(=10\sigma+|z_{\mathrm{obs}}|)$ is sufficient for practical use, where $\sigma$ is the standard deviation of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$. An algorithm for calculating the selective $p$-value via Parametric Programming is summarized in Algorithm 2.
@@
0:  $\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}},z_{\mathrm{min}},z_{\mathrm{max}}$ and $z_{\mathrm{obs}}:=T(\bm{X}_{\mathrm{obs}},\bm{X}^{\mathrm{ref}}_{\mathrm{obs}})$
@@
2:  Obtain $\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$ by (7)
@@
6:     Compute $\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)$ and $\mathcal{M}_{\bm{X}(z)}$ by (26) for $z$
7:     if $\mathcal{M}_{\bm{X}(z)}=\mathcal{M}_{\bm{X}_{\mathrm{obs}}}$ then
8:        $\mathcal{Z}\leftarrow\mathcal{Z}\cup\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)$
@@
10:     $z\leftarrow\max{\mathcal{Z}^{\mathrm{oc}}(\bm{a}+\bm{b}z)}+\gamma$, where $\gamma$ is small positive number.
@@
12:  $p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|Z|>|z_{\mathrm{obs}}|\mid Z\in\mathcal{Z}\right)$
@@ 5.1 Numerical Experiments
Experiments on the type I error rate and power were conducted with two types of covariance matrices: independent $\Sigma=I_{n}\in\mathbb{R}^{n\times n}$ and correlation $\Sigma=(0.5^{|i-j|})_{ij}\in\mathbb{R}^{n\times n}$. In the type I error rate experiments, we used only normal images. The synthetic dataset for normal images is generated to follow $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}\sim\mathcal{N}(\bm{0},\Sigma)$. We made 1000 normal images for $n\in\left\{64,256,1024,4096\right\}$. In the power experiments, we used only abnormal images. We generated 1000 abnormal images $\bm{X}=(X_{1},X_{2},\dots,X_{n})^{\top}\sim\mathcal{N}(\bm{\mu},\Sigma)$. The mean vector $\bm{\mu}$ is defined as $\mu_{i}=\Delta$ for all $i\in\mathcal{S}$, and $\mu_{i}=0$ for all $i\in[n]\backslash\mathcal{S}$, where $\mathcal{S}\subset[n]$ is the anomalous region with its position randomly chosen. The image size of the abnormal images was set to 4096, with signals $\Delta\in\left\{1,2,3,4\right\}$. In all experiments, we made the synthetic dataset for 1000 reference images to follow $\bm{X}^{\mathrm{ref}}=(X^{\mathrm{ref}}_{1},X^{\mathrm{ref}}_{2},\dots,X^{\mathrm{ref}}_{n})^{\top}\sim\mathcal{N}(\bm{0},\Sigma)$. The threshold was set to $\lambda=0.8$, and the kernel size of the averaging filter was set to $3$. All experiments were conducted under the significance level $\alpha=0.05$. The details of the architecture of the diffusion model we used are shown in Appendix C. The diffusion models were trained on the normal images from the synthetic dataset. The diffusion models were trained with $T=1000$ and the initial time step of the reverse process was set to $T^{\prime}=460$, and the reconstruction was conducted $5$ step samplings. The noise schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$ was set to linear. In all experiment, we aim to generate new images through probabilistic sampling, $\eta$ was set to $1$.
In addition, we conducted robustness experiments against non-Gaussian noise. To evaluate the robustness of our proposed method’s performance, we used various non-Gaussian distribution families with different levels of deviation from the standard normal distribution $\mathcal{N}(0,1)$. We considered the following non-Gaussian distributions with a 1-Wasserstein distance $d\in\{0.01,0.02,0.03,0.04\}$ from $\mathcal{N}(0,1)$:
  • Skew normal distribution family (SND).
  • Exponentially modified gaussian distribution family (EMG).
  • Generalized normal distribution family (GND) with a shape parameter $\beta$. This distribution family can be steeper than the normal distribution (i.e., $\beta<2$).
  • Student’s $t$-distribution family ($t$-distribution).
Note that these distributions are standardized in the experiments. Figure 2 shows the probability density functions for distributions from each family, such that the $d$ is set to $0.04$. We run $1000$ trials for each distribution family and each 1-Wasserstein distance to calculate the type I error rate. The significance levels $\alpha$ were set to $0.05$ and $0.10$, and the image size was set to $256$.
Figure 2: Non-Gaussian distributions with $d=0.04$
@@ Results.
Figure 4 shows type I error rate comparison. The proposed methods DMAD-test and DMAD-test-oc can control the type I error rate at the significance level $\alpha$, and bonferroni can control the type I error rate below the significance level $\alpha$. In contrast, naive and permutation cannot control the type I error rate. Figure 4 shows power comparison. Since naive and permutation cannot control the type I error rate, their powers are not considered. Among the methods that can control the type I error rate, the proposed method has the highest power. DMAD-test-oc is over-conditioned and bonferroni is conservative because there are many hypotheses, so they have low power. Figure 5 shows the results of the robustness experiments. DMAD-test maintains good performance on the type I error rate for all the considered distribution families.
@@
(a) SND
(b) EMG
(c) GND
(d) $t$-distribution
Figure 5: Type I Error Rate for Non-Gaussian distribution families
@@ 5.2 Real Data Experiments
We examined the medical image experiment using the dataset extracted from the dataset in Buda et al. (2019). This dataset includes brain images without tumors as normal images (941 images) and brain images with tumors as abnormal images (939 images). A random selection of 100 images from the normal images was made to estimate the mean and variance. The remaining images were standardized using the estimated mean and variance, and 329 normal images were extracted for training the diffusion model. The model was trained with $T=1000$ and the initial time step of the reverse process was set at $T^{\prime}=300$, with reconstruction performed through 5 step samplings. We set the threshold $\lambda=0.9$ and the kernel size of the averaging filter to $3$. The results of the DMAD-test and naive are shown in the Figure 6. The left column represents the results for normal brain images without tumors, while the right column represents the results for abnormal brain images with tumors. The $p_{\rm selective}$ calculated by the proposed DMAD-test is high for normal images (True Negative) and low for abnormal images (True Positive), indicating that the results are desirable. On the other hand, the $p_{\rm naive}$ obtained by the naive test is low not only for abnormal images but also for normal images (False Positive), indicating the invalidness of the naive test. This result indicates that the DMAD-test detected anomalous regions as statistically significant while avoiding misidentification of normal image as anomaly.
$p_{\rm naive}=0.019$, $p_{\rm selective}=0.643$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.379$ Refer to caption $p_{\rm naive}=0.002$, $p_{\rm selective}=0.290$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.822$ Refer to caption $p_{\rm naive}=0.018$, $p_{\rm selective}=0.373$ Refer to caption $p_{\rm naive}=0.024$, $p_{\rm selective}=0.932$ Refer to caption $p_{\rm naive}=0.002$, $p_{\rm selective}=0.225$
(a) Brain images without tumors
$p_{\rm naive}=0.000$, $p_{\rm selective}=0.001$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.002$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.021$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.028$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.027$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.006$ Refer to caption $p_{\rm naive}=0.000$, $p_{\rm selective}=0.016$
(b) Brain images with tumors
@@
Figure 6: An example of the results for applying the proposed DMAD-test and the naive test (an invalid test ignoring that the anomalous region was identified by the diffusion model) to brain images.
6 Conclusions
In this study, we proposed a statistical test for anomalous regions in medical images detected by a diffusion model. With the proposed DMAD-test, the false detection rate can be controlled with the significance level because statistical inference is conducted conditional on the fact that the anomalous regions are identified by a diffusion model. We believe this study marks a step toward bridging the gap between generative AI and rigorous statistical inference in medical imaging analysis.